OpenStudy (anonymous):

Evaluate cot -220 degrees. round answer to the nearest hundredth

OpenStudy (anonymous):

Good morning, my child.\[\cot(-220^{\circ}) = \dfrac{\cos(-220^{\circ})}{\sin(-220^{\circ})} \]

hartnn (hartnn):

or since cot has period of 180, \(\cot (-220) = \cot (-220+180)=\cot (-40)\) or even, \(\cot(-40)=\cot(-40+180)=\cot(140)\)